If joy can paint a box alone in 6 hours and Gina can paint it in 4 hours. How many hours would it take both to paint the box together?
Answer by TimothyLamb(4379)

r = w/t
rates are additive:
1/6 + 1/4 = 0.41666666666
---
t = w/r
1/0.41666666666 = 2.4 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php